# Vaidya-Reissner-Nordström Extension On the White-hole Region and the White Hole Evaporation Dynamic

### v1

@@ Abstract
We present an analytic construction of a radiating charged white hole solution by Vaidya-Reissner–Nordström metrics. And find the classic Vaidya-Reissner–Nordström white hole evaporation process dynamic. Beginning with the maximal analytic extension of the classical Schwarzschild white hole via Kruskal–Szekeres coordinates, we establish a foundation to explore the intricate causal structure of white hole spacetime. Incorporating the Vaidya metric in double-null coordinate, we introduces a dynamic element to the model through a monotonically decreasing mass function, representing the effects of outgoing null dust and radiative mass loss. By extending this framework to include electric charge, we derive the Vaidya–Reissner–Nordström metric, where both mass and charge decrease over retarded time, figure out the influence of electromagnetic fields on the evolution of the white hole.
@@ Introduction
Our analysis reveals that the Vaidya–Reissner–Nordström metric not only generalizes the static Reissner–Nordström geometry but also provides an effective framework for modeling the evaporation process of white holes. The evolution of the Bondi mass and charge under outgoing radiation is shown to preserve the weak energy condition, thereby ensuring consistency with cosmic censorship.
@@ I Schwarzschild white hole
For the start doing with the white hole with the Vaidya-Reissner-Nordström white hole (actually the charged, non-rotation white hole), we need to start with the most basic white hole, the Schwarzschild solution which maximal extended. We have the advanced (ingoing) and retarded (outgoing) situation. And as the white hole is just the time-reverse for the black hole, and non one can entry the ”event horizon” (actually it is anti-horizon). For the simplest solution is asymptotically flat solution of Schwarzschild, can be express in the Schwarzschild coordinate (In the paper we apply the geometrodynamic units, G=c=1):
@@
The Schwarzschild metric have two singularities, $r=0$ and $r=2M$. $r=2M$ is the coordinate-set singularity, and $r=0$ is the physical singularity. To avoid the coordinate singularity, we need to enroll the Kruskal–Szekeres coordinate. First we apply the Tortoise coordinate transfer[2], $r^{\star}=2M\log{(\frac{r}{2M}-1)}$ and $dr^{\star}=\frac{dr}{1-\frac{2M}{r}}$, This coordinate stretches the radial coordinate so that the horizon $r=2M$ been sent to $\infty$- ∞ in $r^{\star}$, and the null coordinate we apply:
@@
The transformation will clash when $r=2M$ (curvature singularity) in Eq.(9) and Eq.(10).
@@
The Penrose-Carter diagram of the white hole region for the Schwarzschild spacetime shown as figure 2:
Figure 2: Penrose diagram
@@ II Vaidya white hole
with $\epsilon=1$ for advanced (ingoing) time and $\epsilon=-1$ for retarded(outgoing) time. Since we aim to reach the non-static property of the Vaidya white hole, we note that the Kruskal-Szekeres and Eddington-Finkelstein coordinate do not behave well in the time-reversed dynamic region. More detailed, although the Kruskal-Szekeres coordinate system is a form of double-null coordinates, it was specifically designed to extend the Schwarzschild metric, and the Eddington-Finkelstein coordinates are less effective in studying the causal structure and dynamic evolution of spacetime in a symmetric null framework. Therefore, we apply a general double-null coordinate system. In our study of the white hole, we use retarded time $(\epsilon=-1$) because of the mass is outgoing. And the mass function $M(\omega)$ is monotonically decreasing, meaning $\dot{M}(\omega)\leq 0$ ($\dot{M}=\frac{\partial M}{\partial t}$)[4]. This choice of double-null coordinates allows us to construct the analytic maximal extension for the Vaidya white hole region in a manner that is both more intuitive and physically reasonable for the time-reversed scenario. We construct the double-null coordinate as the form:
@@
where $d\Omega^{2}=d\theta^{2}+\sin^{2}\theta d\phi^{2}$ is the unit 2-sphere. Here $r(u,v)$ is the areal radius, and $f(u,v)$ is an undetermined metric function to be solved for in the next illustration. The metric signature is (,+,+,+)(-,+,+,+)( - , + , + , + ) provided $f>0$. The coordinates $u$ and $v$ are chosen to be null, so surfaces of constant $u$ or $v$ are null hypersurfaces.
For an outgoing radiation field, we take $\epsilon=-1$, meaning the null dust flows along the $u$-direction (outgoing) and the mass decreases with increasing $u$ (retarded time). The stress-energy tensor is that of a single directional null dust:
@@
with $k_{a}$ a null vector field tangent to outgoing radial null geodesics (so $k_{a}\propto\nabla_{a}u$). This $T_{ab}$ has only a $uu$-component, consistent with purely outgoing radiation. Physically, $\rho(u)$ is related to the mass loss rate. In particular, we assume a mass function $m(u)$ such that $M(u)$ is monotonically decreasing in $u$. The Einstein equations then imply $T_{uu}=\rho(u),(k_{u})^{2}=G_{uu}$, which gives the standard Vaidya relation for mass loss:
$\frac{dM(u)}{du}=-4\pi r^{2}T_{u}^{u}<0,$ (14)
so that $T_{uu}=\frac{-M^{\prime}(u)}{4\pi r^{2}}$ (with $M^{\prime}=dM/du$) is the energy density of the outgoing null flux. All other $T_{ab}$ components vanish, and $M(u)$ will turn out to be the Bondi mass as measured at null infinity. Given the above ansatz, we now write down the nontrivial Einstein field equations. Spherical symmetry and the null dust form of $T_{ab}$ yield a system of equations for $f(u,v)$ and $r(u,v)$. In particular, the mixed $uv$-component (the trace of the field equations) gives a vanishing Ricci scalar. From $T_{a}^{a}{}=0$ we have $R_{a}^{a}{}=0$, which leads to:
$r_{uu}^{{}^{\prime}}=\frac{f_{u}^{{}^{\prime}}}{f}r_{u}^{{}^{\prime}}$ (15)
where $r_{u}^{{}^{\prime}}=\partial r/\partial u$.
$02\frac{r_{u}^{{}^{\prime}}r_{v}^{{}^{\prime}}+rr_{uv}^{{}^{\prime}}}{f}+1=$ 0 (16)
Substituting the equation (15) into this, we obtain an conservation law:
$2\frac{r_{u}^{{}^{\prime}}r_{v}^{{}^{\prime}}}{f}=\frac{2M(u,v)}{r}-1,$ (17)
where we have defined a local mass function $M(u,v)$ by $M(u,v)=\frac{r_{u}^{{}^{\prime}}r_{v}^{{}^{\prime}}r}{f}+\frac{r}{2}$ followed by Eq.(13). In fact, the above can be recognized as the defining equation for the Misner–Sharp mass in spherical symmetry. Rearranging, one convenient definition is:
$1-\frac{2M(u,v)}{r}\equiv-\frac{2r_{u}^{{}^{\prime}}r_{v}^{{}^{\prime}}}{f}$ (18)
@@
Here $B(u)$ is an arbitrary function of $u$ arising from integration. Equation (19) is precisely the single first-order nonlinear PDE that the Einstein equations reduce to for the Vaidya metric. It governs how the areal radius $r(u,v)$ evolves as one moves along increasing $v$, for each fixed $u$. The function $B(u)$ reflects the residual freedom to rescale the null coordinates. In other words, $B(u)$ defines the scale of the $u$-coordinate; one can exploit this freedom to simplify the metric or impose convenient normalization conditions. The weak energy condition requires $M^{\prime}(u)\leq 0$ for an outgoing solution, and one can choose $B(u)$ positive so that $u$ increases toward the future.
Since $M(u)$ is fixed with respect to $v$, this integration Eq.(20) yields an implicit solution for $r(u,v)$:
$r+2M(u)\ln|r-2M(u)|=B(u)v+C(u).$ (21)
Here $C(u)$ is an arbitrary integration function of $u$. (The logarithmic term arises from the integral of $dr/(1-2M/r)$, as in the Schwarzschild-Kruskal solution in the section I.) The function $C(u)$ corresponds to a further freedom in the null coordinate $v$. A suitable choice of $C(u)$ will be made to ensure the coordinates are regular across any horizon; we discuss this shortly. Now, with $r(u,v)$, we can construct the metric function $f(u,v)$. One convenient way to obtain $f$ is to use the relation $g^{uv}=-1/f$ together with the definition of $M(u)$. Equivalently, we can differentiate the implicit solution above with respect to $v$. Differentiating yields back the radial PDE (19), $r_{v}^{{}^{\prime}}=B(u)\big{(}1-2M/r\big{)}$. Thus, solving for $f=-2g_{uv}$:
$f(u,v)=2B(u)\frac{\partial r}{\partial v}=2B(u)B(u)\left(1-\frac{2M(u)}{r}\right),$ (22)
@@
$f(u,v)=2B^{2}(u)\left(1-\frac{2M(u)}{r(u,v)}\right).$ (23)
This is the desired metric coefficient in double-null coordinates, valid for an arbitrary mass function $M(u)$. We emphasize that $B(u)$ is an arbitrary gauge function which reflecting the freedom to redefine $u$, and $C(u)$ is another free function coming from the integration of Equation(19).
To make contact with familiar forms, one may exploit the coordinate freedom to simplify $B(u)$ and $C(u)$. For example, in asymptotically flat cases one convenient choice is to set $u$ to coincide with retarded time at infinity, which fixes $B(u)$ in the limit of large $r$. One natural condition is $B(u)\to 1$ as $r\to\infty$ so that $u$ measures proper time at infinity. Likewise, $C(u)$ can be chosen to eliminate spurious coordinate singularities: for instance, one demands that the metric be analytic across the outgoing horizon (the surface $r=2m(u)$). This regularization condition picks out a unique $C(u)$ that extends the metric through $r=2m$ analogous to the Kruskal extension for Schwarzschild. In the static case $m(u)=M$ (constant), setting $B(u)=1$ and $C(u)=0$ yields $f(u,v)=2(1-2M/r)$ and the implicit solution reduces to $r+2M\ln|r-2M|=v$, which are precisely advanced Eddington–Finkelstein coordinates. A different choice of $C(u)$ would produce the standard Kruskal–Szekeres coordinates (maximally extended, regular at the horizon). In general, choosing $C(u)$ appropriately allows one to cover the outgoing Vaidya spacetime smoothly across any event horizon or Cauchy horizon that forms.
By taking $\epsilon=-1$ (outgoing radiation) at the outset, we have derived the white-hole version of the Vaidya metric. The line element can now be written in its final form. Dropping the explicit $(u,v)$ dependence, the metric is:
$ds^{2}=-2f(u,v)dudv+r^{2}(u,v)d\Omega^{2},$ (24)
$f(u,v)=2B^{2}(u)\left(1-\frac{2M(u)}{r(u,v)}\right).$
Together with the implicit relation for $r(u,v)$, this describes a general radiating white-hole spacetime. The outgoing Vaidya (white-hole) metric is often presented in retarded null coordinates $(u,R)$ (a single-null coordinate form). To recover that form, one can use the freedom in $v$ to set $C(u)$ such that $v$ coincides with the ingoing null coordinate of flat infinity. Then $u$ is our retarded time, and $R=r$ is the areal radius. The metric becomes:
$ds^{2}=-\left(1-\frac{2M(u)}{R}\right)du^{2}-2dudR+R^{2}d\Omega^{2}.$ (25)
@@
In summary, we started from Einstein’s equations with a null dust stress-energy and arrived at a single first-order PDE for $r(u,v)$. Its integration introduced an arbitrary function of $u$, which we denoted $C(u)$, that must be specified by a regularity condition. The general solution for arbitrary $M(u)$ is given implicitly by the relations above. The metric function $f(u,v)$ is determined explicitly in closed-form by $M(u)$ and the gauge choice $B(u)$. For the white-hole case, $m(u)$ is a decreasing function (outgoing positive energy flux), and one picks the integration constants to produce an outgoing radiation solution regular on the outgoing horizon. Our construction here follows the formalism of Waugh and Lake[5], generalizing the Schwarzschild-Kruskal coordinates to the dynamic case of a changing mass. The result is a double-null Vaidya metric.
@@ III Vaidya white hole evaporation
III.1 Vaidya White Hole and Double-Null Formalism
@@
where $\sigma(u,v)$ is a conformal factor which arbitrary function for $u$ and $v$. $u$ (ingoing) and $v$ (outgoing) are independent null coordinates. The Vaidya solution can transfer in this form, but it is often simpler to use one of the Eddington–Finkelstein forms above with an explicit mass function. In these coordinates, Einstein’s equations can be solved given a specified $M(u)$ or $M(v)$, and the stress-energy will correspond to a null dust flow. Below we focus on the outgoing Vaidya metric for an evaporating white hole with $M=M(u)$ decreasing.
@@
For the stress-energy to be physically sensible as positive energy flux leaving the object, we require $M^{\prime}(u)<0$ so that $\rho>0$. Indeed, the null energy condition $T_{ab}k^{a}k^{b}\geq 0$ for all null $k^{a}$ demands $M^{\prime}(u)\leq 0$, a negative slope indicates the central mass is decreasing as it radiates energy outward. This models an evaporating white hole.
Should note that, in the ingoing Vaidya metric (advanced time $v$), one finds instead $R_{vv}=\frac{2M^{\prime}(v)}{r^{2}}$ and $T_{ab}=\frac{M^{\prime}(v)}{4\pi r^{2}}n_{a}n_{b}$ with $n_{a}=-\partial_{a}v$. Then $M^{\prime}(v)>0$ is required by energy conditions, corresponding to increasing mass due to infalling radiation (e.g. a collapsing star or accreting black hole). In our radiating white hole case , the outgoing formulation with $M^{\prime}(u)<0$ is used. Notably, allowing $M^{\prime}(u)<0$ or $M^{\prime}(v)<0$ formally violates the usual energy conditions. Since one needs effectively negative energy influx to reduce mass. This is a classical surrogate for Hawking radiation, which indeed involves violation of classical energy conditions. The Vaidya solution itself is still a solution of Einstein’s equations, but the stress-energy for a truly evaporating hole is that of a null dust with negative energy density falling in, or equivalently positive energy streaming out to infinity. (However, the violation for the energy condition is not impossible, experimentally shown in Casimir effect.[6][7][8][9][10])
@@
III.2 Mass Loss and Outgoing Energy Flux

### v2

@@ Abstract
We develop an analytic model that extends classical white hole geometry by incorporating both radiative dynamics and electric charge. Starting from a maximal analytic extension of the Schwarzschild white hole via Kruskal Szekeres coordinates, we introduce a time dependent mass function, representative of outgoing null dust to model evaporation. Building on this foundation, the study then integrates the Reissner-Nordström framework to obtain a dynamic, charged white hole solution in double null coordinates. In the resulting Vaidya Reissner Nordström metric, both the Bondi mass and the associated charge decrease monotonically with retarded time, capturing the interplay of radiation and electromagnetic effects. Detailed analysis of horizon behavior reveals how mass loss and charge shedding modify the causal structure, ensuring that energy conditions are preserved and cosmic censorship is maintained.
@@ Introduction
Our analysis reveals that the Vaidya–Reissner–Nordström metric not only generalizes the static Reissner–Nordström geometry but also provides an effective framework for modeling the evaporation process of white holes. We generalized Hiscock evaporation black hole model in the white hole situation. The evolution of the Bondi mass and charge under outgoing radiation is shown to preserve the weak energy condition, thereby ensuring consistency with cosmic censorship.
@@ I Schwarzschild white hole
We need to start with a basic Schwarzschild white hole, by apply the Schwarzschild solution which with maximal extended. The white hole is the time-reverse edition for the black hole, The metric can be express as:
@@
The Schwarzschild metric have two singularities, $r=0$ and $r=2M$. $r=2M$ is the coordinate-set singularity, and $r=0$ is the true singularity. To avoid the coordinate singularity, we need to enroll the Kruskal–Szekeres coordinate. First we apply the tortoise coordinate[2], $r^{\star}=r+2M\ln{|\frac{r}{2M}-1|}$ and $dr^{\star}=\frac{dr}{1-\frac{2M}{r}}$, This coordinate stretches the radial coordinate so that the horizon $r=2M$ been sent to $\infty$- ∞ in $r^{\star}$, and the null coordinate we apply:
@@
The transformation will clash when $r=2M$ for curvature singularity in Eq.(9) and Eq.(10).
@@
The Penrose-Carter diagram of the white hole region for the Schwarzschild spacetime shown as figure 2.
@@ II Vaidya white hole
with $\epsilon=1$ for advanced time and $\epsilon=-1$ for retarded time.
Since we aim to reach the non-static property of the Vaidya white hole, we note that the Kruskal-Szekeres and Eddington-Finkelstein coordinate do not behave well in the time-reversed dynamic region.
More detailed, although the Kruskal-Szekeres coordinate system is a form of double-null coordinates, it was specifically designed to extend the Schwarzschild metric, and the Eddington-Finkelstein coordinates are less effective in studying the causal structure and dynamic evolution of spacetime in a symmetric null framework.
Therefore, we apply a general double-null coordinate system. In our study of the white hole, we use retarded time because of the mass is outgoing. And the mass function $M(\omega)$ is monotonically decreasing, $\dot{M}=\frac{\partial M(\omega)}{\partial t}\leq 0$[4]. The construction in the double-null coordinate as the form[5]:
@@
where $d\Omega^{2}=d\theta^{2}+\sin^{2}\theta d\phi^{2}$ is the unit 2-sphere, $r(u,v)$ is the areal radius, and $f(u,v)$ is an undetermined metric function to be solved for later.
Given $f>0$ to remain the metric signature be (,+,+,+)(-,+,+,+)( - , + , + , + ). The coordinates $u$ and $v$ are chosen to be null, so surfaces of constant $u$ or $v$ are null hypersurfaces.
For an outgoing radiation field, we take $\epsilon=-1$, meaning the null dust flows along the $u$-direction and the mass decreases with increasing retarded time $u$. The stress-energy tensor is that of a single directional null dust:
@@
with $k_{a}$ a null vector field tangent to outgoing radial null geodesics and we have $k_{a}\propto\nabla_{a}u$. We take $k_{a}=\nabla_{a}u$, by set the normalization $k_{u}=1$ and all other $k_{v},k_{\theta},k_{\phi}=0$. Hence $T_{uu}=\rho(u)$.
$T_{ab}$ only has a $uu$-component, consistent with purely outgoing radiation. Physically, $\rho(u)\geq 0$ is related to the mass loss rate. In particular, we assume a mass function $M(u)$ is monotonically decreasing in $u$. The standard Vaidya relation for mass loss:
$\frac{dM(u)}{du}=-4\pi r^{2}T_{uu}<0,$ (14)
so that $T_{uu}=\frac{-M^{\prime}(u)}{4\pi r^{2}}$ (with $M^{\prime}=dM/du$) is the energy density of the outgoing null flux.[6] All other $T_{ab}$ components vanish, and $M(u)$ will turn out to be the Bondi mass as measured at null infinity.
Given the above ansatz, we now write down the nontrivial Einstein field equations. Spherical symmetry and the null dust form of $T_{ab}$ yield a system of equations for $f(u,v)$ and $r(u,v)$. In vacuum situation $T_{uu}=0$, we have:
$r_{uu}=\frac{f_{u}}{f}r_{u}$ (15)
where noted $r_{u}=\partial r/\partial u$, $r_{v}=\partial r/\partial v$. The Einstein equations with outgoing null dust $T_{ab}=\rho(u)\nabla_{a}u\nabla_{b}u$ given:
$G_{uu}=-\frac{2}{f}\left(r_{uu}-\frac{f_{u}}{f}r_{u}\right)=8\pi T_{uu}\Longrightarrow r_{uu}-\frac{f_{u}}{f}r_{u}=-4\pi fT_{uu}(u)$
$G_{uv}=-\frac{2\left(r_{u}r_{v}+rr_{uv}\right)}{fr}+\frac{f}{2}=0\Longrightarrow 2\frac{r_{u}r_{v}+rr_{uv}}{f}=f\frac{r}{2}.$
And we have:
$2\frac{r_{u}r_{v}}{f}=\frac{2M(u,v)}{r}-1,$ (16)
where we have defined a local mass function $M(u,v)=\frac{r_{u}r_{v}r}{f}+\frac{r}{2}$. In fact, the above can be recognized as the defining equation for the Misner–Sharp mass in spherical symmetry. Rearranging, one convenient definition is[7]:
$1-\frac{2M(u,v)}{r}\equiv-\frac{2r_{u}r_{v}}{f}$ (17)
@@
Here $B(u)$ is an arbitrary function of $u$ arising from integration.
Equation (18) is precisely the single first-order nonlinear PDE that the Einstein equations reduce to for the Vaidya metric. It governs how the areal radius $r(u,v)$ evolves as one moves along increasing $v$, for each fixed $u$. $B(u)$ can be chosen positive so that $v$ increases toward the future for each fixed $u$. Performing the $r$-integral in Eq.(19) one obtains:
$r-2M(u)+2M(u)\ln|r-2M(u)|=B(u)v+C(u),$ (20)
Here $C(u)$ is an arbitrary integration function of $u$, where the additive constant $-2M(u)$ may be absorbed into the arbitrary function $C(u)$. A suitable choice of $C(u)$ will be made to ensure the coordinates are regular across any horizon.
Now, with $r(u,v)$, we can construct the metric function $f(u,v)$. One convenient way to obtain $f$ is to use the relation $g^{uv}=-1/f$ together with the definition of $M(u)$. Equivalently, we can differentiate the implicit solution above with respect to $v$. Differentiating yields back the radial PDE (19), $r_{v}=B(u)\big{(}1-2M/r\big{)}$. Thus, solving for $f=-g_{uv}$:
$f(u,v)=\frac{r_{v}}{B(u)},$ (21)
@@
$f(u,v)=1-\frac{2M(u)}{r(u,v)}.$ (22)
This is the desired metric coefficient in double-null coordinates, valid for an arbitrary mass function $M(u)$. $B(u)$ gauges the normalization of $v$ along each outgoing null cone with fixed $u$. $C(u)$ gauges the origin of $v$ on each cone. Neither of them changes the scaling of the retarded coordinate $u$.
One may now exploit the coordinate freedom encoded in the functions $B(u)$ and $C(u)$. First, fix the overall scaling of the ingoing coordinate $v$ by choosing $B(u)=1$. Which makes the retarded coordinate $u$ coincide with the usual Bondi retarded time at future null infinity. Second, use the additive freedom $v\rightarrow v+\sigma(u)$ to select $C(u)$ so that the metric remains analytic at the outgoing apparent (or event) horizon $r=2M(u)$; in the static Schwarzschild case this gives the familiar Kruskal choice $C(u)=-2M$.
With those choices the implicit solution of the radial equation is:
$r+2M(u)\ln|r-2M(u)|=v,$
and the metric coefficients are:
$ds^{2}=-2\left(1-\frac{2M(u)}{r}\right)dudv+r^{2}d\Omega^{2},$ (23)
$f(u,v)=1-\frac{2M(u)}{r(u,v)}.$
In the static limit $M(u)=M$ these reduce to the advanced Eddington-Finkelstein form; with a different $C(u)$ one recovers the Kruskal-Szekeres extension. For a dynamical $M(u)$, a suitably chosen $C(u)$ likewise ensures a smooth extension across any future (white-hole) horizon that forms.
Together with the implicit relation for $r(u,v)$ in Eq.(20), this describes a general radiating white-hole spacetime. The outgoing Vaidya white hole metric is often presented in retarded null coordinates $(u,R)$. To recover that form, one can use the freedom in $v$ to set $C(u)$ such that $v$ coincides with the ingoing null coordinate of flat infinity. Then $u$ is our retarded time, and $R=r$ is the areal radius. The metric becomes:
$ds^{2}=-\left(1-\frac{2M(u)}{R}\right)du^{2}-2dudR+R^{2}d\Omega^{2},$ (24)
@@
In summary, we started from Einstein’s equations with a null dust stress-energy and arrived at a single first-order PDE for $r(u,v)$. Its integration introduced an arbitrary function of $u$, which we denoted $C(u)$, that must be specified by a regularity condition. The general solution for arbitrary $M(u)$ is given implicitly by the relations above. For the white-hole case, $M(u)$ is a decreasing function (outgoing positive energy flux), and one picks the integration constants to produce an outgoing radiation solution regular on the outgoing horizon. Our construction generalizes the Schwarzschild-Kruskal coordinates to the dynamic case of a changing mass. The result is a double-null Vaidya metric[5].
@@ III Vaidya white hole evaporation
III.1 Basic Structure
@@
where $\sigma(u,v)$ is a conformal factor which arbitrary function for $u$ and $v$. $u$ and $v$ are independent null coordinates. The Vaidya solution can transfer in this form, but it is often simpler to use one of the Eddington–Finkelstein forms above with an explicit mass function. In these coordinates, Einstein’s equations can be solved given a specified $M(u)$ or $M(v)$, and the stress-energy will correspond to a null dust flow. Below we focus on the outgoing Vaidya metric for an evaporating white hole with $M=M(u)$ decreasing.
@@
For the stress-energy to represent positive energy flux leaving the object, we take
$\rho=-\frac{M^{\prime}(u)}{4\pi r^{2}}>0\quad\Longrightarrow\quad M^{\prime}(u)<0.$
The null-energy condition
$T_{ab}k^{a}k^{b}\geq 0$
with $T_{ab}=-\frac{M^{\prime}(u)}{4\pi r^{2}}l_{a}l_{b}$ therefore requires $M^{\prime}(u)\leq 0$; for the evaporating case $M^{\prime}(u)<0$ and the NEC is satisfied. In the ingoing Vaidya metric one instead has
$T_{ab}=\frac{M^{\prime}(v)}{4\pi r^{2}}n_{a}n_{b}$
Here the NEC implies $M^{\prime}(v)\geq 0$, describing accretion or collapse. Forcing the mass to decrease in this chart $\left(M^{\prime}(v)<0\right)$ would indeed require negative energy density and violate the classical energy conditions an idealized stand-in for quantum effects such as Hawking radiation, much as the Casimir vacuum violates point wise energy conditions. And the violation for the energy condition is not impossible, experimentally shown in Casimir effect by [8][9][10][11][12].
@@
III.2 Mass Loss and Outgoing Energy Flux with Hiscock Model